OpenStudy (anonymous):

Fan + Medal. Geoff planted dahlias in his garden. Dahlias have bulbs that divide and reproduce underground. In the first year, Geoff's garden produced 8 bulbs. In the second year, it produced 16 bulbs, and in the third year it produced 32 bulbs. If this pattern continues, how many bulbs should Geoff expect in the 6th year? A) 64 bulbs B) 512 bulbs C) 128 bulbs D) 256 bulbs My Answer: D) 256 bulbs I found my answer by adding like this, 8 + 8 = 16, then 16 + 16 = 32 and then 32 + 32 = 64, 64 + 64 = 128, and then 128 + 128 to get 256. Please check! Thanks.

OpenStudy (mrnood):

Yes - but it would be 'normal to consider it as 8*2=16 16*2= 32 32*2 = 64 etc. i.e. the number is doubling each year. Your answer is correct and your method is OK, but if you are about to start looking at 'progressions' and series then seeing it as x2 each year will be more helpful in future..

OpenStudy (anonymous):

Alright, thank you. @MrNood